Fermi National Laboratory


Accelerator Update

Monday July 5

The day shift began with Operations monitoring the Recycler stash and the Antiproton Source (Pbar) stack. There was no Linac beam due to a problem with an RF station (LRF5).

At 11:24 AM, a Linac expert reported that LRF5 was having a water problem. Operators investigated further and discovered a temperature regulation problem. They made adjustments to warm the RF cavity to its nominal range.

Another Linac expert arrived to help with LRF5 at 11:54 AM.

Linac experts brought the LRF5 RF station back online. Operations established beam to the Linac dump at 1:56 PM.

Pbar resumed stacking at 1:58 PM.

Operations established beam to MiniBooNE at 2:14 PM.

Operations put the Tevatron (TeV) into shot setup at 2:48 PM.

Operations established store 3619 at 4:55 PM with an initial luminosity of 60.02E30.

At 5:09 PM, Operations reported that a collimator hadn't moved to its proper position. The Collider Detector at Fermilab (CDF) turned off their detector and complained about colliding beam problems. Operators called a TeV expert.

At 5:54 PM, a Main Injector (MI) RF station (MIRF15) tripped off. Operators investigated and decided at 6:27 PM that they could run without this station and bypassed it.

Operations halted stacking at 6:40 PM and began stashing antiprotons into the Recycler.

Operations reported at 6:56 PM that a TeV D0 cryo compressor tripped off. Cryo system technicians at CHL bypassed the compressor.

A Booster RF station (BRF16) tripped off at 7:31 PM and would not reset. Operators bypassed the station and called in an expert.

Pbar resumed stacking at 7:41 PM.

A Pbar expert began a three-hour zero-stack study around 8 PM.

A Booster expert replaced a control unit for BRF16 and had it back on line by 8:57 PM.

A Booster RF station (BRF12) tripped off at 9:05 PM. It reset okay.

Operations established beam to the Meson MCenter experiment at 10:55 PM.

Tuesday July 6

The midnight shift began with Operations monitoring store 3619 and the stash, with Pbar stacking, and with beam to MiniBooNE and MCenter experiments taking beam.

At 1:19 AM, a Booster RF station (BRF1) tripped off. It reset okay.

At 2:15 AM, BRF12 tripped off for the first of many times over the next three hours. It was eventually bypassed.

At 6:07 AM, BRF4 tripped off and wouldn't reset. Operators investigated and then called in an expert who reported at 6:40 AM that he was having trouble getting the station on. Operations had to lower the Booster beam intensity until at 7:05 AM when they managed to coax BRF12 to turn on.

Operations began establishing beam to the Meson MTest experiment at 7:04 AM.

BRF12 tripped off at 7:15 AM. It reset okay.

At 7:54 AM, a Pbar Debuncher power supply (D:H704) tripped off. Operators contacted an expert to investigate.

BRF12 tripped off at 8:05 AM. It reset okay.

A Pbar expert reported at 8:37 AM, that D:H704 was fixed and back online.

Stacking resumed at 8:41 AM.

At 8:52 AM, the MiniBooNE beam safety system permit tripped off due to a power supply (E:H866) not staying within its current limits. Operators contacted experts.

BRF12 tripped off at 9:01 AM. It reset okay.

Operation disabled beam for MCenter at 12:04 PM.

Techs investigating the MiniBooNE power supply E:H866 replaced a voltage regulator and a broken connector at 12:16 PM. Operations reestablished beam to MiniBooNE.

At 1:35 PM, Operations turned off all the MI RF stations to allow techs to repair MIRF15. They had the station repaired with Pbar stacking and MTest taking beam by 1:56 PM.

RF techs brought BRF4 back on line at 3:02 PM.

Switchyard (SY) experts began, at 6 PM, a two-hour SY dump study.

Operations terminated store 3619 at 7 PM.

Cryo system techs began wet engine work at TeV sector A3 and D2 service buildings at 7:17 PM. They completed their work thirty minutes later.

CHL brought the D0 cryo compressor back on line at 7:28 PM.

Operations put the TeV through a wet squeeze at 8:07 PM.

Operations put the TeV into a Mixed Antiproton Mode shot setup at 8:36 PM.

After experts dealt with a Pbar cooling problem, a MI coalescing problem, and a Recycler transfer problem, Operations established store 3621 at 11:23 PM. Its initial luminosity was 92.1E30. This set a NEW RUN II RECORD LUMINOSITY. Pbar resumed stacking.

Operations reported at 11:25 PM that they had to make some collimator adjustments to resolve a colliding beam problem.

A TeV RF station (TRF8) tripped off at 11:40 PM.

Wednesday July 7

The midnight shift began with Operation monitoring store 3621, with Pbar stacking, and with MiniBooNE and MTest taking beam.

Operations began sending beam to MCenter around 12:06 AM.

BR12 tripped off twice at 2:11 AM. It tripped off three more times over the next three hours.

At 6:05 AM, the TeV suffered a sector D1 quench possibly due to a vacuum problem that caused a cryo heat load.

Operations prepared the TeV, MI, Booster, SY, and MiniBooNE for access around 7 AM.

Operations prepared CDF and D0 for a three-hour access at 8:11 AM.

Vacuum techs reported at 8:11 AM that they discovered ice-covered kautzky valve in sector D1.

Plans for Wednesday Morning

The plan for today is to fix the D1 kautzky valve, repair BRF12, install a new antiproton source target, and conduct maintenance in MI and other areas. All accesses except for Pbar should be completed by noon. The Pbar target work should be completed by late this afternoon. The next TeV shot will come sometime this evening.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/7/2004   email Fermilab